meriah:

Trip walks a total of 12 miles to and from school each week. Lindsey rides her bike a total of 56 miles to and from school each week. Use the greatest common factor and the distributive property to complete the equation, and find the total number of miles Trip and Lindsey travel to and from school each week. (3 + ) =

justus:

12 + 56 GCF = 4 (4 * 3) + (4 * 14) = 4(3 + 14) <==